# Effect of realistic oscillator phase noise on the performance of cell-free networks

### v1

@@ Abstract
To keep supporting 6G requirements, the radio access infrastructure will increasingly densify. Cell-free (CF) networks offer extreme flexibility by coherently serving users with multiple Access points (APs). This paradigm requires precise and stable phase synchronization. In this article, we adapt the standardized 5G NR setup (subcarrier spacing, OFDM symbol duration and allocation) to investigate the effect of Phase Noise (PN) on the simulated performance of scalable CF networks. In contrast to the prior literature relying on the simplified model of a free-running oscillator with the Wiener process, we deploy a realistic hardware-inspired phase noise model reproducing the Local Oscillator (LO) phase drift. Our results demonstrate that even affordable LOs offer sufficient stability to ensure negligible loss of uplink Spectral Efficiency (SE) on the time scale of the standardized 5G Transmission Time Interval of 1 ms. This study substantiates the feasibility of CF networks based on 5G standards.
Index Terms:
6G, cell-free, phase synchronization, MIMO
@@ I Introduction
Cell-free (CF) networking will play a pivotal role in the upcoming 6G networks [1, 2]. In CF, user devices (often called ”User Equipment” or UE) are served by multiple Access Points (APs) performing coherent joint transmission and reception [3]. This approach offers superior Spectral Efficiency (SE) compared to conventional architectures like Macro Base Stations or small cells [4], promising a more efficient and effective network infrastructure for 6G.
@@
CF networking has attracted attention for its impressive performance in theoretical studies under ideal conditions [4]. Building upon this success, subsequent research has focused on incorporating more realistic assumptions, including the utilization of non-ideal hardware, to facilitate practical implementation. These efforts can be categorized into two main groups: 1) hardware implementations or testbeds aimed at validating and demonstrating the feasibility of CF networking in real-world scenarios, and 2) analytical and simulation studies that employ more realistic assumptions. Since the main enabler of the CF is coherent processing, here we would like to underline the importance of phase synchronization between the communication nodes (e.g., a UE and APs serving the user, including inter-AP synchronization).
Testbeds
@@
Wang et al. [5] implemented a centralized cloud-based CF mMIMO network. In their testbed, a CPU served 16 eight-antenna UEs with 16 eight-antenna APs. In [6], the KU Leuven CF Multiple-Input-Multiple-Output (MIMO) testbed was used to obtain channel state information (CSI) in a dense CF deployment consisting of 8 APs with 8 antennas each (64 antennas in total). In both works, most of the Hardware Impairments (HWI) effects are implicitly assessed since real equipment was used. Unfortunately, both works used a common clock and the effect of Phase Noise (PN) was not estimated.
Theoretical studies
Several works investigated the effect of Local Oscillator (LO) PN on a single carrier [7, 8] and on OFDM-based [9] CF networks. The articles assess the performance of CF networks when non-ideal LOs are used. Inspired by Petrovic et al. [10], the authors of the papers modeled PN by a discrete-time independent Wiener process assuming independent innovations at every time instance. The papers reported significant degradation in comparison to an ideal LO: up to 58% performance degradation in terms of 5%-outage rate in [8]. The authors of [7, 9] propose channel estimators compensating for PN in the sense of estimating the effective channel. By utilizing the assumed structure of PN in the estimator and then using the estimate for receive combining, a sort of phase noise compensating signal processing framework was proposed.
Stat-of-the-Art limitations: The PN model used in [7, 8, 9] is a good choice for a first approximation of real LO performance. However, considering the immense importance of coherent processing for the potential of CF, more realistic PN models must be used to assess the CF network performance as it was highlighted by practitioners [11].
Contributions: In this work, we
  1. 1.
@@
    Critically analyze the impact of time-frequency resource allocation specified by 3GPP on PN.
  2. 2.
@@
    Deploy a realistic hardware-inspired phase noise model reproducing the LO phase drift.
  3. 3.
@@
    Simulate a CF network to assess the effect of realistic PN on uplink SE of 3GPP-compliant OFDM system.
Notation: Vectors and matrices are denoted by boldface lower and upper case symbols. The symbols $(\cdot)^{T}$, $(\cdot)^{H}$, and $(\cdot)^{\dagger}$ express the transpose, Hermitian transpose, and pseudo-inverse operators, respectively. The notations $\mathbb{C}^{L\times 1}$ and $\mathbb{C}^{W\times W}$ refer to complex $L$-dimensional vectors and $W\times W$ matrices, respectively. Finally, $\mathbf{a}\sim\mathcal{CN}(0,\mathbf{\Sigma})$ represents a circularly symmetric complex Gaussian variable with zero mean and covariance matrix $\mathbf{\Sigma}$.
Figure 1: CF massive MIMO network with user-centric clustering.
@@ II System Model
First, we introduce the OFDM system model considering PN effects. Secondly, we adopt standard 3GPP settings and show their effect on the expected influence of PN. Finally, we introduce the scalable CF network model taking into account the realistic effects of PN on the OFDM system.
Industry-baked efforts targeting the development of CF networks are heavily based on 5G NR standards [12, 13] to simplify the transition to practice. In this article, we use the same approach. To focus solely on the communication performance, we assume that the Initial Access, Pilot Assignment, Cluster Formation, Channel Estimation, and Radio Resource Allocation procedures are completed by utilizing techniques available in the literature [14, 15].
II-A OFDM system model
Consider an OFDM system with $N$ subcarriers, a subcarrier spacing $\Delta f$, and a bandwidth $B=1/T_{\rm s}=N\Delta f$, where $T_{s}$ is the sampling period. Signal $\{X_{i}\}_{i=0}^{N-1}$ is transmitted over $N$ subcarriers of an OFDM symbol, with an average per-symbol power $p$. By means of an $N$-point IDFT, we can write the time-domain representation of the OFDM symbol transmitted by UE as
@@ II-A OFDM system model
where index $n$ is the time-domain channel use (sample).
II-A1 Phase Noise Model
Time-dependent random phase drifts, also known as PN, cause multiplicative distortions of the received signal. The distortions are introduced when the signal is multiplied by the LO’s output during up- and down-conversion operations.
The total PN between one user and one AP at $n$-th time sample can be expressed as
$\theta_{n}=\phi_{n}+\varphi_{n}.$ (2)
The particular realizations of the $\phi_{n}$ and $\varphi_{n}$ depend on the model, e.g. it can be a simple Wiener process or a random process with a complex spectrum defined by various instabilities of the LOs (phase noise, phase random walk, frequency noise, frequency random walk etc.). The models of the phase noise are discussed in details in Sec. III.
We can express time-domain PN realization during one received OFDM symbol as $\mathbf{p}^{\rm{time}}=[e^{j\theta_{0}},\dots,e^{j\theta_{N-1}}]^{T}\in\mathbb{C}^{N\times 1}$ while its frequency-domain counterpart is written as $\mathbf{p}^{\rm{freq}}=[P_{0},\dots,P_{N-1}]^{T}\in\mathbb{C}^{N\times 1}$ where the elements are obtained via DFT as $P_{i}=\frac{1}{N}\sum_{n=0}^{N-1}e^{j\theta_{n}}e^{-j2\pi ni/N}$.
II-A2 OFDM Signal Model with Phase Noise
@@
PN effects on the received signal for each subcarrier $i\in\{0,1,\cdots,N-1\}$, (4) can be rewritten as [16]
@@
where $(\cdot)_{N}$ denotes the modulo-$N$ operation. The Common Phase Error (CPE) is an identical phase rotation in all subcarriers within an OFDM symbol while the Inter-Carrier Interference (ICI) is an additive noise (not always Gaussian). Note that for a given OFDM symbol, CPE can be estimated and compensated during the channel estimation process.
II-A3 Channel Model
We consider a time-varying channel that is divided into Coherence Blocks (CBs) with a coherence time $T_{\rm c}$ and a coherence bandwidth $B_{\rm c}$. Within the coherence time, $N_{\rm ct}\triangleq\lfloor{T_{\rm c}}/{T_{\rm OFDM}}\rfloor$ OFDM symbols with duration $T_{\rm OFDM}$ are transmitted. Similarly, we can calculate the number of subcarriers within the coherence bandwidth as $N_{\rm cb}\triangleq\lfloor{B_{\rm c}}/{\Delta f}\rfloor$. Each CB spans $N_{\rm ct}$ and $N_{\rm cb}$ successive OFDM symbols and subcarriers, over which the channel is time-invariant and frequency flat. In the following, let us consider one coherence block as the focus of our analysis.
Numerology: For the Frequency Range 1 (FR1: sub-6 GHz), the most popular 5G NR numerology is $\mu=0$, consequently, the smallest time-frequency resource unit that can be scheduled to a user corresponds to a CB spanning over $B_{\rm c}$=180 kHz bandwidth ($N_{\rm cb}=12$ subcarriers with $\Delta f=15$ kHz spacing) and $T_{\rm c}=1$ ms (divided into $N_{\rm ct}=14$ OFDM symbols with a duration of $T_{\rm OFDM}=$71.4 $\mu$s each). The number of coherent channel uses is $\tau=N_{\rm ct}N_{\rm cb}$.
@@ II-A4 Resource Element Mapping and Channel Estimation
In line with [15], we assume the Transmission Time Interval (TTI) format where all OFDM symbols are used for uplink communication. To enable channel estimation and CPE compensation, at least one OFDM symbol should contain pilot signals (Demodulation Reference Signal) according to 3GPP TS 38.211. Though most commonly OFDM symbols 3 or 4 are used, we can assume a scenario where the pilots are transmitted at the very beginning of the TTI (see Fig. 2). The coherent channel uses $\tau=\tau_{p}$ and $\tau_{c}$ consist of channel uses for pilot and data, respectively.
@@
The pilot allocation in Fig. 2 represents the worst-case scenario from the CPE point of view. The CPE is compensated only at $\mathfrak{t}=0$ OFDM symbol while the other symbols are affected by the growing error. For an arbitrary OFDM symbol $\mathfrak{t}$, the received signal is expressed as:
@@
where $P_{0}^{\mathfrak{t}}$ is the CPE for the OFDM symbol $\mathfrak{t}$ (note that $P_{0}^{0}=1$) and $H$ is constant over $N_{\rm ct}$ OFDM symbols and $N_{\rm cb}$ subcarriers. Compared to (5), this equation also takes into account that the 5G subcarrier spacing aims at eliminating the effect of ICI.
II-B Scalable Cell-Free network
The CF massive MIMO (CF mMIMO) network shown in Fig. 1 is a communication network consisting of $M$ APs distributed over a geographical area that collaborate to serve $K$ users through coherent joint transmission and reception [3] utilizing shared time-frequency resources. The APs are connected to central processing units (CPUs) responsible for coordination. Precoding and combining operations can be performed at each AP, centrally at the CPU, or through a distributed approach involving both APs and CPUs [6]. In this study, we focus on the centralized processing.
II-B1 CF Channel model
Let us consider a general case of APs equipped with $L$ antennas and extend the model in Sec II-A3. Within a coherence block, the random channel vector between UE $k$ and AP $m$ is modeled by independent correlated Rayleigh fading as $\mathbf{h}_{mk}\sim\mathcal{CN}(0,\mathbf{R}_{mk})$ where the small scale fading is modeled with complex Gaussian distribution and $\mathbf{R}_{mk}\in\mathbb{C}^{L\times L}$ is a deterministic positive semi-definite correlation matrix that describes the large-scale propagation effects (i.e., shadowing and pathloss) with a fading coefficient $\beta_{mk}=tr(\mathbf{R}_{mk})/L$. Similarly to [4, 7], we assume that the set of correlation matrices is known by the communication nodes.
@@ II-B2 Scalable CF
To ensure network scalability, each AP serves just a set of UEs. Björnsson and Sanguinetti [4] suggested forming clusters of APs serving users in a user-centric manner (see Fig. 1). In this case, AP $m$ serves just a subset of $K$ UEs ensuring the network scalability
$\mathcal{D}_{m}=\{k:\quad\mathrm{tr}(\mathbf{D}_{mk})\geq 1,\quad k\in\{1,\dots,K\}\}.$ (7)
II-B3 Phase Noise in CF
Since there is no ICI, phase noise should be updated once per OFDM symbol $\mathfrak{t}$ which corresponds to an update of CPE every $T_{\rm OFDM}$. For a practical scenario where $L$ antennas of the $m$-th AP are connected to a single LO, the corresponding PN matrix for $k$-th UE is expressed as $\mathbf{\Theta}_{mk}^{\mathfrak{t}}=e^{j\theta_{mk}^{\mathfrak{t}}}\mathbf{I}_{L}$, where $\theta_{mk}^{\mathfrak{t}}$ is obtained as in (2) for AP $m$ and UE $k$. Note that $\mathbf{\Theta}_{mk}^{0}=\mathbf{I}_{L}$ as CPE for $\mathfrak{t}=0$ is included in the channel estimation.
II-B4 Received Signal Model
Here we derive the achievable SE for a practical scalable CF network with PN. We use centralized processing, where the APs forward all received signals to the CPU. Next, the CPU performs the channel estimation and data detection. Note that to isolate the negative effects of channel estimation error from the effects introduced by PN, we assume perfect channel estimation.
In uplink, the $M$ APs send the received signals $\{\mathbf{y}_{m,i}:m=1,\cdots,M\}$ to the CPU. In such case by setting $W=ML$, the uplink signal received on $i$-th subcarrier of the given ODFM symbol can be written as

### v2

# Effect of Realistic Oscillator Phase Noise on the Performance of Cell-Free Massive MIMO Systems

@@ Abstract
As the demand for 6G technologies continues to grow, the radio access infrastructure is expected to become increasingly dense. Cell-free (CF) Massive MIMO systems provide remarkable flexibility by enabling coherent service to users through multiple Access Points (APs). This innovative paradigm necessitates precise and stable phase synchronization. This paper examines the standardized 5G New Radio (NR) framework, focusing on subcarrier spacing, OFDM symbol duration, and allocation, while investigating the impact of Phase Noise (PN) on the performance of scalable massive MIMO cell-free systems. Unlike existing studies that typically employ a simplified model of a free-running oscillator characterized by a Wiener process, we present a realistic phase noise model inspired by actual hardware, designed to accurately capture the Local Oscillator (LO) phase drift. Furthermore, our PN model extends its applicability beyond cell-free systems, making it relevant for any RF system operating within the sub-6 GHz band. This model provides a robust foundation for the practical design of cell-free systems, encompassing numerology and pilot allocation strategies. Our findings reveal that even cost-effective low-cost Local Oscillators can achieve sufficient stability, resulting in negligible degradation of uplink Spectral Efficiency (SE) within the standardized 5G Transmission Time Interval of 1 ms. These results affirm the viability of cell-free massive MIMO systems based on 5G standards and their potential integration into future 6G networks.
@@ I Introduction
Cell-free (CF) massive multiple input-multiple output (mMIMO) systems are poised to play a critical role in the evolution of 6G networks [1, 2] as they address targets such as a 10-fold increase in user density, spectral efficiency (SE), and data rate compared to 5G. In CF mMIMO, user devices, commonly referred to as User Equipment (UE), are served by multiple Access Points (APs) that perform coherent joint transmission and reception [3]. This innovative approach offers significantly enhanced SE compared to traditional architectures, such as Macro Base Stations or small cells [4], thereby promising a more efficient and effective network infrastructure for the forthcoming 6G era.
@@
CF mMIMO has garnered considerable attention due to its remarkable performance demonstrated in theoretical studies conducted under ideal conditions [4]. Building on this foundation, subsequent research has aimed to incorporate more realistic assumptions, including the use of non-ideal hardware, to facilitate practical implementation. These efforts can be classified into two primary categories: 1) hardware implementations or testbeds designed to validate and demonstrate the feasibility of CF mMIMO systems in real-world scenarios, and 2) analytical and simulation studies that adopt more realistic assumptions. Since coherent processing is the primary enabler of CF mMIMO, the phase synchronization between communication nodes (e.g., a UE and the APs serving it) is of paramount importance.
@@
a) Testbeds: Wang et al. [5] implemented a centralized cloud-based CF mMIMO system, where a CPU served 16 eight-antenna UEs with 16 eight-antenna APs. In another study [6], the KU Leuven CF Multiple-Input-Multiple-Output (MIMO) testbed was utilized to obtain channel state information (CSI) in a dense CF mMIMO deployment comprising 8 APs, each equipped with 8 antennas (64 antennas in total). In both studies, the effects of Hardware Impairments (HWI) were implicitly assessed due to the use of real equipment. However, both works relied on a common clock, and the impact of Phase Noise (PN) was not evaluated.
b) Theoretical Studies: Several investigations have explored the influence of Local Oscillator (LO) PN on single carrier [7, 8] and OFDM-based CF mMIMO systems [9]. Inspired by [10], papers [7, 8, 9] modeled PN by a discrete-time independent Wiener process, assuming independent innovations at each time instance. They observed significant performance degradation compared to ideal LO conditions, with drops of up to 58% in terms of the 5%-outage rate [8] and a tenfold decrease in Spectral Efficiency within a 1 ms communication interval [9]. The authors of [7, 9] proposed channel estimators that compensate for PN by estimating the effective channel. Specifically, they leveraged the assumed structure of PN in the channel estimator, subsequently using the PN-corrected channel estimates to calculate the Maximum Ratio (MR) and Minimum Mean-Square Error (MMSE) combining vectors, which are then employed to combine signals received from multiple APs into a single enhanced signal.
The PN model utilized in [7, 8, 9] offers a reasonable approximation of actual LO performance. However, in [7, 8], the LO phase drifts with a variance of 0.23 radians per OFDM symbol (71.4 us), whereas hardware manufacturers report phase noise that is an order of magnitude smaller for reasonably comparable frequency and time scales in affordable Software Defined Radios [11]. This discrepancy highlights the need for an improved PN model.
The contributions of this paper are as follows:
@@
  • We provide a critical analysis of the impact of 3GPP time-frequency resource allocation on PN.
@@
  • We introduce a realistic, hardware-inspired phase noise model that accurately reproduces LO phase drift.
@@
  • We simulate a CF mMIMO system to evaluate the effect of realistic PN on the uplink SE of a 3GPP-compliant OFDM system.
To isolate and accurately assess the impact of PN, our study intentionally excludes other factors such as channel estimation errors and channel aging from our model.
@@ II System Model
In this section, we introduce the OFDM system model considering the effects of PN. Next, we adopt standard 3GPP settings to analyze their impact on the anticipated influence of PN. Finally, we present a scalable OFDM CF mMIMO system model, reflecting the realistic effects of PN.
Industry efforts in developing CF mMIMO systems are largely based on the 5G New Radio (NR) standards [12, 13], which facilitates a smoother transition to practical implementation. In this paper, we follow a similar approach. To concentrate solely on communication performance, we assume that the procedures for Initial Access, Pilot Assignment, Cluster Formation, Channel Estimation, and Radio Resource Allocation are completed using techniques documented in the literature [14, 15].
II-A The OFDM System Model
Consider an OFDM system comprising N subcarriers, with a subcarrier spacing of $\Delta f$ and a bandwidth defined as $B=1/T_{\rm s}=N\Delta f$ where $T_{s}$ represents the sampling period. The transmitted signal $\{X_{i}\}_{i=0}^{N-1}$ is conveyed over $N$ subcarriers of an OFDM symbol, with an average per-symbol power of $p$. Utilizing an $N$-point Inverse Discrete Fourier Transform (IDFT), we can express the time-domain representation of the OFDM symbol transmitted by the UE as follows, where the index $n$ denotes the time-domain channel use (sample).
@@ II-A The OFDM System Model
II-A1 The Phase Noise Model
Time-dependent random phase drifts, commonly known as PN, induce multiplicative distortions in the received signal. These distortions arise when the signal is multiplied by the output of the LO during the up- and down-conversion processes.
The total PN experienced between a user and an AP at the $n$-th time sample can be expressed as follows
$\theta_{n}=\phi_{n}+\varphi_{n},$ (2)
where $\phi_{n}$ and $\varphi_{n}$ represent PN at the LOs of UE and AP, respectively. The particular realizations of the $\phi_{n}$ and $\varphi_{n}$ depend on the model (see Sec. III).
We can express the time-domain PN realization during one received OFDM symbol as $\mathbf{p}^{\rm{time}}=[e^{j\theta_{0}},\dots,e^{j\theta_{N-1}}]^{T}\in\mathbb{C}^{N\times 1}$ while its frequency-domain counterpart is written as $\mathbf{p}^{\rm{freq}}=[P_{0},\dots,P_{N-1}]^{T}\in\mathbb{C}^{N\times 1}$ where the elements are obtained by DFT as $P_{i}=\frac{1}{N}\sum_{n=0}^{N-1}e^{j\theta_{n}}e^{-j2\pi ni/N}$.
II-A2 The OFDM Signal Model with Phase Noise
@@
The PN affected frequency domain received signal $Y$ for each subcarrier $i\in\{0,1,\cdots,N-1\}$ are expressed as [16]
@@
where $(\cdot)_{N}$ denotes the modulo-$N$ operation.
The Common Phase Error (CPE) $P_{0}$ represents an identical phase rotation across all subcarriers within an OFDM symbol, while the additive noise $P_{\ell}$ (which is not always Gaussian) constitutes the Inter-Carrier Interference (ICI) between carriers $i$ a and $\ell$. It is important to note that for a given OFDM symbol, the CPE can be estimated and compensated for during the channel estimation process.
II-A3 The Channel Model
We consider channel segmented into Coherence Blocks (CBs) characterized by a coherence time $T_{\rm c}$ and a coherence bandwidth $B_{\rm c}$. Within the coherence time, $N_{\rm ct}\triangleq\lfloor{T_{\rm c}}/{T_{\rm OFDM}}\rfloor$ OFDM symbols, each with a duration of $T_{\rm OFDM}$, are transmitted. Similarly, we can calculate the number of subcarriers within the coherence bandwidth as $N_{\rm cb}\triangleq\lfloor{B_{\rm c}}/{\Delta f}\rfloor$. Each CB encompasses $N_{\rm ct}$ and $N_{\rm cb}$ successive OFDM symbols and subcarriers, during which the channel remains time-invariant and frequency flat. For our analysis, we will focus on one CB.
Numerology: For Frequency Range 1 (FR1: sub-6 GHz), the predominant 5G NR numerology is $\mu=0$. Consequently, the smallest time-frequency resource unit allocated to a user corresponds to a CB spanning a bandwidth of $B_{\rm c}$=180 kHz ($N_{\rm cb}=12$ subcarriers with $\Delta f=15$ kHz spacing and a coherence time of $T_{\rm c}=1$ ms (comprising $N_{\rm ct}=14$ OFDM symbols, each with a duration of $T_{\rm OFDM}=$71.4 $\mu$s each). The total number of coherent channel uses is $\tau=N_{\rm ct}N_{\rm cb}$.
@@ II-A4 Resource Element Mapping and Channel Estimation
In accordance with [15], we assume a Transmission Time Interval (TTI) format where all OFDM symbols are utilized for uplink communication. To facilitate channel estimation and CPE compensation, at least one OFDM symbol must include pilot signals (Demodulation Reference Signal) as per 3GPP TS 38.211. While OFDM symbols 3 or 4 are most commonly employed, we can consider a scenario where pilots are transmitted at the very beginning of the TTI (see Fig. 1. The coherent channel uses $\tau=\tau_{p}$ and $\tau_{c}$ comprise channel uses for pilot and data, respectively.
@@
The pilot allocation depicted in Fig. 1 represents a worst-case scenario from the CPE perspective. The CPE is compensated only during the $\mathfrak{t}=0$ OFDM symbol, while subsequent symbols are affected by an increasing error. For an arbitrary OFDM symbol $\mathfrak{t}$, the received signal can be expressed as follows.
@@
where $P_{0}^{\mathfrak{t}}$ is the CPE for the OFDM symbol $\mathfrak{t}$ (note that $P_{0}^{0}=1$) and $H$ is constant over $N_{\rm ct}$ OFDM symbols and $N_{\rm cb}$ subcarriers. Similarly, $X_{i}^{\mathfrak{t}}$ and $Z_{i}^{\mathfrak{t}}$ respectively denote the transmit symbol and noise for the subcarrier $i$ of the OFDM symbol $\mathfrak{t}$.
Compared to (5), this equation also accounts for the fact that the 5G subcarrier spacing is designed to mitigate the effects of ICI.
II-B The Scalable Cell-Free system
The CF mMIMO system depicted in Fig. 2 represents a communication network comprising multiple APs strategically distributed across a geographical area. These APs collaborate to serve $K$ users through coherent joint transmission and reception, utilizing shared time-frequency resources [3]. Each AP is connected to CPUs responsible for coordinating the network. The operations of precoding and combining can occur at individual APs, centrally at the CPU, or through a hybrid approach involving both APs and CPUs [6]. In this paper, we focus on centralized processing.
II-B1 The CF Channel Model
We consider a general scenario where APs are equipped with $L$ antennas, extending the model presented in Sec II-A3. Within a coherence block, the random channel vector between UE $k$ and AP $m$ is modeled as independent correlated Rayleigh fading, represented by $\mathbf{h}_{mk}\sim\mathcal{CN}(0,\mathbf{R}_{mk})$ where, the small-scale fading is characterized by a complex Gaussian distribution, and $\mathbf{R}_{mk}\in\mathbb{C}^{L\times L}$ denotes a deterministic positive semi-definite correlation matrix that captures the large-scale propagation effects, such as shadowing and path loss, with a fading coefficient defined as $\beta_{mk}=tr(\mathbf{R}_{mk})/L$. Consistent with [4, 7], we assume that the set of correlation matrices is known to the communication nodes.
@@ II-B2 Scalable CF
Figure 2: CF massive MIMO system with user-centric clustering.
To promote network scalability, each AP is assigned to serve a specific set of UEs. Björnsson and Sanguinetti [4] proposed the formation of clusters of APs that serve users in a user-centric manner (see Fig. 2). In this configuration, AP $m$ exclusively serves a subset $\mathcal{D}_{m}$ of $K$ UEs, ensuring network scalability, defined as follows:
$\mathcal{D}_{m}=\{k:\quad\mathrm{tr}(\mathbf{D}_{mk})\geq 1,\quad k\in\{1,\dots,K\}\},$ (7)
where $\mathbf{D}_{mk}\in\mathbb{C}^{L\times L}$ is a diagonal matrix indicating the communication links between APs and UEs.
II-B3 The Phase Noise in CF
Given the absence of ICI, the phase noise is updated once per OFDM symbol, corresponding to an update of the CPE every $T_{\rm OFDM}$. In practical scenarios, where the $L$ antennas of the $m$-th AP are connected to a single LO, the phase noise matrix for the $k$-th UE can be expressed as $\mathbf{\Theta}_{mk}^{\mathfrak{t}}=e^{j\theta_{mk}^{\mathfrak{t}}}\mathbf{I}_{L}$ where $\theta_{mk}^{\mathfrak{t}}$ is derived as indicated in (2) for AP $m$ and UE $k$. It is important to note that $\mathbf{\Theta}_{mk}^{0}=\mathbf{I}_{L}$ as the CPE for $\mathfrak{t}=0$ is included in the channel estimation.
II-B4 The Received Signal Model
In this section, we derive the achievable SE for a practical scalable CF mMIMO system with phase noise. As previously mentioned, we consider centralized processing, where all received signals from the APs are forwarded to the CPU. The CPU then performs channel estimation and data detection. To isolate the negative effects of channel estimation errors from those introduced by phase noise, we assume perfect channel estimation. In the uplink scenario, the $M$ APs transmit the received signals $\{\mathbf{y}_{m,i}:m=1,\cdots,M\}$ to the CPU.
By setting $W=ML$, the uplink signal received on the $i$-th subcarrier of the given OFDM symbol can be expressed as: